27 September 2011 UT: MODS checkout

02:11 Active Preset sent. Collimating. Stars show lots of astigmatism, coma and some trefoil. Takes about 7 minutes to reach WFE 400-500 nm.

02:19 Send StoneO_1520 active preset. Dual imaging, 60 sec. g and r imaging started at 02:24. Was observing manually --- locked IMCS manually... Blue image was not transferred. We checked in CRB (>MODS1.BC PING). Still not saving. Need to restart data transfer server for Blue. Blue 2 and Red 4 are StoneO_1520. Seeing estimated to be 0.8" on blue (Blue camera has pixel scale: 0.120 arcsec/pixel) and 0.63" (Red camera has pixel scale: 0.123 arcsec/pixel).

02:43 Two 120 sec exposures on StoneO_1520 field at g and r. Two more 120s exposures at u and i. i image shows parasitic light (common to both channels but stronger in red than blue). image is number 6. Two more 120s at u and z (number 7). z also shows parasitic light.

02:53 Closed hatch and taking g and i for 120 seconds to see whether we still see the light leak. No parasitic light when hatch closed.

02:55 Active Preset to StoneO_1477 (field with more stars than 1520). Guide star seeing simultaneous with images with 0.8/-.63 arc was 1.0". Guide stage focus is a little soft. The different guide cameras have different depths and while focussed in lab, may not be in focus. There is not consistecy between the depths of the guide cameras. Will be corrected once MODS off telescope.

02:58 g and r 120s images. On axis guide star. Probe in field. A set at g (0.8") and i (0.6"). Parasitic light with fringing on i image. u (180 sec) and z (120 sec). Guider showing 0.9" seeing.

03:30 acqMODS sdss2013. Red only. 13 slit image 14 field image. 15 is confirmatory thru-slit image. 16-19 are 250s spectra (0.6" slit, red only, gg495 filter) of sdss2013. (note added later: I must have meant sdss2103).

04:05 LDS749B standard star. 30 sec wth same config. 19 is slit image, 20 field image. modsAlign2 -l (19) (20). 21 is confirmatory thru-slit. Got script error due to bug in execMODS script (Rick id's it and fixed it now). =ERROR: Command timed out after 60 seconds. Abort, Retry, Ignore? We checked IMCS lock (restarted it from obs4 with modsCmd) and then waited for it to complete and finally hit "i" for Ignore. Script ended but set of images continue. (If this happens again, we will hit "retry". If, after the timeout period of 120 seconds, it still does not lock, there is a problem)

04:32 acqMODS pks2216.acq. modsAlign -l (26) (27). 28 is confirmatory.

04:39 execMODS pks2216.obs. Dual spectra: 0.6" slit. No blocking on Blue and GG495 blocking on Red. 900sec per channel. Note, this uses 2 900sec, but recommended to do 3 images for CR rejection. CR on thick CCDs significant, in particular in red

use acq images from the night to gauge seeing (since focus is soft on guider)

05:16 acqMODS bd284211.acq (standard star). Preset failed --- way out of collimation (astigmatism and out of focus). Was at 82 deg elevation. Going instead to Feige110.

05:25 modsAlign -l (31) (32) on Feige110. 33 is confirmatory.

05:27 execMODS feige110.obs Red seeing on acq image 0.67" and guider gave 0.93".

05:39 acqMODS wolfrayet.acq. Uses 1" slit and is at about 60 deg elevation. modsAlign -l 38,39 40 is confirmatory. When star comes up through-slit, need to mark slit at a position different, slightly above, since the "x" on the slit makes a fit through the slit. When star is saturated, you see bleeding and depressed counts in center.

05:50 execMODS wolfrayet.obs. GCS star is bright. Put in ND filter, but still reporting bright. We see GS elongated @ low WFS RMS and guide cam exp time below lower limit. The GS is not consistently round - shape is changing and red + is bouncing around (within blue circle). Bright star at 60 deg elevation. Guide star is far off axis. A row of lenslets is not illum and therefore this is interpreted as trefoil.

05:57 Noticed Blue fits images were not transferred from 10 to 17... Rick fixed this and the blue images started to appear.

06:16 acqMODS feige110.acq (for ND wolfrayet program). Ran this while reading out (which is the way to go as far as efficiency - has been extensively tested and there is no resultant noise on the image). But script stalled at PARTNER... eventually continued. 49 and 50 are slit and field images. 51 confirmatory image. About 2800 counts. [Should aim for higher - 20k counts. Usually 1 or 2 (cosmic ray rejection will be OK with this for standards) 180 or 300s exposures are recommended].
  • In command line on interface:
    • exptime 180
    • nimgs 1
    • go

06:43 acqMODS sn2006gy.acq 1" slit. red 57 and 58. Can use modsdisp display to look into scaling/details.

06:53 execMODS sn2006gy.obs I sent this during the readout of the last acq exposure, but got into a race condition. An error was trapped: Error: nimgs. Cannot change this parameter during image acquisition. Abort, Retry,Ignore? I entered Retry. The error was trapped and it went along OK.

00:25 acqMODS OSU_m33obXmods.acq =ERROR: PRESET IIF error - PresetTelescope result status: Error, Stop[65] complete, couldn't find a star in image. Abort, Retry or Ignore? = David corrected pointing and I then hit "r" to retry. Working now. Elevation 77 deg.

0042 execMODS OSU_m33obXmods.obs started. 4 1200 sec exposure. Not on parallactic, but high. 21.6 mag, 16.5 mag guide star. 1:20 - 33 msec motion (AGW test report).

08:27 Suddenly we lost the guide star. We were not tracking well at this elevation (87.8 deg elevation). Object rose from 77 deg to 87deg. Fortunately we lost the guide star just after the 2nd of 4 exposures ended and was reading out.

08:38 Copied the script to Xmods2.obs and edited it to have nimgs 2 (vs original nimgs 4). Running acquisition. 68 and 69 are slit and field images. 70 confirmatory. On modsdisp display we see an image filename like:
mods1r.20110927.0068.fits  - OBJECT - M33 Object X Acquisition - r_sdss - LS5x60x0.6 Mask - IN
110923M8.04b.fits   - OBJECT - M33 Object X Acquisition - r_sdss - LS5x60x0.6 Mask - IN
mods1r.20110927.0069.fits  - OBJECT - M33 Object X Acquisition - r_sdss - LS5x60x0.6 Mask - STOW
DO NOT TRY TO OPEN a file like this because we do not know its origin. For some reason, Red sometimes writes two copies of the same image. It uses this type of "unique filename" so as not to overwrite the first copy.

08:50 Seeing on last acq image was 0.5". The guider is reading about 0.9".

09:09 Clouds encroaching. Just as 4th spectrum was started. I typed pause in the Command window. The pause behavior was not expected. Instead of pausing, closing the shutter and keeping the integration countdown fixed, it closed the shutter and readout the very short integrations (images 35 and 72, blue and red). Then the script was reported to be done. Pause acted more like a stop.

09:46 Taking a pair of Blue/Red images in g/i filters to check for light leak. Dome is dark and shut. Telescope pointing at horizon. Nothing seen in the images.

09:57 Running calibrations at zenith. Dome lights off. Maybe dome lights can be on during lamps, but not during biases or flat fields. "gr" grating. "bgr" blue only grating. "rgr" red only grating.
  • execMODS grlamps.cal - 1 of each lamp in blue and red: Ne+Hg[Ar], Xe+Kr, Ar (no Ne in blue and no Hg in red). (note: only one lamp flat needed for MOS)
  • execMODS grslitflat.cal
  • execMODS grflats.cal (0.6" slit) Quart Lamp Spectral Flat. (made a local copy of script in /home/OSURC/MODS/Calib/)
  • execMODS grflats1arcsec.cal (1.0" slit). Used same exposure times as for 0.6". Gets up to 56k in Blue with both qth1 and qth2.

11:19 execMODS modsSleep.pro FITSFLUSH done as part of this script. Note that many Red images had not been transferred.

11:22 doBias (in ~MODSeng/modScripts/Calib/) which takes 10 spec biases (8K x 3K) and 10 imaging biases (3K x 3K). Readout time is 72 sec, so 20 x 72 sec --- will take about 25 minutes to complete.

General Notes about MODS
  • Blue spectrum, +x --- +lambda. Red spectrum +x --- -lambda
  • Only time when you can send a script during readout is when it is an acq following an obs script
  • 5577 OI line auroral line on blue spec at X=6410pix. 5577 OI on red spec at X= 6710 pix. Dichroic cuts off at 5650 angstroms
  • Blue Grating spectrum centered at 4410 angstroms
  • Red Grating spectrum centered at 7730 angstroms

-- OlgaKuhn - 27 Sep 2011
Topic revision: r7 - 26 Nov 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback